OpenStudy (meganhall0517):

Which equation has a solution of -5? a. 2a – 6 = 4 b. 3a + 7 = 8 c. 3a – 7 = 2 d. 3a + 19 = 16 4 5

OpenStudy (brainzonly):

A equals positive 5... B equals 1/3... C equals 3... D equals -5 D

OpenStudy (meganhall0517):

Which equation represents the following sentence? One-forth of a number plus five equals the number minus seven. a. 1/4n + 7 = n – 5 b. 1/4n + 5 = n – 7 c. 4n + 7 = n – 5 d. 4n + 5 = n – 7

OpenStudy (brainzonly):

@meganhall0517 B. think about it! 1/4 of a number (n) 1/4n Plus 5: 1/4n + 5 equals the number minus 7 1/4n + 5 = n - 7